OpenStudy (anonymous):

Suppose you won a special math lottery and were given the following options: Option A- You are given $10,000 and recieve $1,000 every year thereafter. Option B- You are given $1,000 and recieve 10% more each year. Which option would you choose and why? What if the lottery went bankrupt in five years, are the plans equal? Would you take a one-time payout of $50,000 instead of option A or B? Why or why not?

OpenStudy (anonymous):

b becasue i would spend all the money at one time :D

OpenStudy (anonymous):

in 5 years A is still better than B and it depends how long you expect to live

OpenStudy (anonymous):

Can someone please answer the bankrupt question though? I don't get how I would solve that...

OpenStudy (anonymous):

A=10000 + 4(1000) B=1000 + 1100 + 1210 + 1331 + 1464.1 = 6105.1

OpenStudy (anonymous):

So after that's all figured out, option A would still be better than a $50,000 one-time payout?

OpenStudy (anonymous):

If you are going to live 40 or more years (ignoring interest altogether) A is better but I think $50,000 is better because you can invest it and earn interest and if you dont live more than 40 years (probably more than 45-50) you make out better